In a batch of 15 jars of instant coffee, 5 have been under filled. Suppose 6 of these jars are selected without replacement. If more than 1 of these selected jars are under filled, the whole batch is sent back for refill. What is the probability of the batch being sent back for refill?



1
Expert's answer
2022-05-04T14:40:00-0400

Let "X=" the number of  jars under filled: "X\sim Bin(n, p)."

Given "n=6, p=5\/15=1\/3, q=1-p=2\/3."


"P(X>1)=1-P(X=0-P(X=1)"

"=1-\dbinom{6}{0}(\dfrac{1}{3})^0(\dfrac{2}{3})^{6-0}-\dbinom{6}{1}(\dfrac{1}{3})^1(\dfrac{2}{3})^{6-1}"

"\approx0.65"


The probability of the batch being sent back for refill is 0.65.